# EntGPT: Linking Generative Large Language Models with Knowledge Bases

### v1

@@ 4.2. Comparison Models
  • End2End (Kolitsas et al., 2018): The End-to-End Neural Entity Linking (End2End) is a global ED and entity linking system based on Deep-ED (Ganea and Hofmann, 2017), but with a more robust recurrent neural network (RNN) architecture that considers mention-context and mention-mention interactions.
  • REL (van Hulst et al., 2020): The Radboud Entity Linker (REL) is an ED and entity linking package which combines the NER method FLAIR (Akbik et al., 2019) and ED method mulrel-nel (Le and Titov, 2018).
  • GENRE (Cao et al., 2021): The Generative Entity REtrieval model (GENRE) is a generation-based model that re-casts the ED task as an entity name generation process.
  • ner4el (Tedeschi et al., 2021): The Named Entity Recognition for Entity Linking (NER4EL) model uses named entity recognition to perform the entity linking task, which is similar to the entity disambiguation task.
  • ReFinED (Ayoola et al., 2022b): The Representation and Fine-grained typing for Entity Disambiguation (ReFinED) model is an efficient end-to-end zero-shot-capable EL model is a recent system that considers Wikipedia entries as extra features to the ED and entity linking tasks.
  • ExtEnd (Barba et al., 2022b): The Extractive Entity Disambiguation (ExtEnd) model is a transformer-based framework where the disambiguation task is recast as a machine reading comprehension task.
  • CoherentED (Xiao et al., 2023b): The Coherent Entity Disambiguation (CoherentED) model aims to enhance the coherence of entity predictions and achieves state-of-the-art performance on most datasets.
@@
  • ROBERTa (Liu et al., 2019): The Robustly Optimized BERT Pretraining Approach (ROBERTa) is an optimized version of the Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) model, which uses Masked Language Modeling (MLM) as its pre-training objective. The pre-trained model is fine-tuned on each question answering benchmark in a fully supervised manner.
@@
  • QAGNN & GreaseLM (Yasunaga et al., 2021; Zhang et al., 2022a): The Question Answering Graph Neural Network (QAGNN) and the Graph REASoning Enhanced Language Model (GreaseLM) both fine-tune an LLM using a knowledge graph (KG) on downstream tasks. These architectures are broadly referred to as KG-augmented question answering models.
@@
  • DRAGON (Yasunaga et al., 2022): Deep Bidirectional Language-Knowledge Graph Pretraining (DRAGON) uses a knowledge base to enhance the pre-training stage and also performs fine-tuning on downstream tasks in a supervised manner.
@@
In the zero-shot setting, we use the following LLMs as foundation models: Large Language Model Meta AI (Llama, Llama2, & Llama2-chat(Touvron et al., 2023a, b), the Fine-tuned LAnguage Net (FLAN-137B) (Wei et al., 2021), and the Generative Pre-trained Transformer (GPT3.5) (Ouyang et al., 2022).
Table 1. Data statistics for 11 experimental datasets. Number of documents (# of Docs), number of mentions (# of Mention).
AIDA MSN AQU ACE04 CWEB WIKI KORE OKE15 OKE16 REU RSS
# of Docs 231 20 50 35 320 319 50 101 173 113 357
# of Mention 4485 656 727 257 11154 6793 144 536 288 650 524
4.3. Evaluation Metrics
@@
To maintain a fair comparison across datasets, we use the in-KB micro-F1 score as our evaluation metric for ED following the example of Guo and Barbosa (2018). Specifically, being in-KB requires that ground truth mentions correspond to existing knowledge base entries. Empty or out-of-KB entities are removed in the evaluation process. The mean Micro-F1 score is taken per-mention. Although a model may predict non-entities, each mention will have some corresponding ground truth entity. For question answering, we use accuracy as the evaluation metric.
4.4. Results and Findings
@@
4.4.1. Entity Disambiguation (RQ1 and RQ2)
Table 2. Test micro-F1 scores on standard six benchmarks. With each setting, best scores are highlighted in bold. † indicates that the results are obtained directly from the source paper.
Fine-tuning Data Model AIDA MSN AQU ACE04 CWEB WIKI avg
Wiki+AIDA REL (van Hulst et al., 2020) 0.928 0.935 0.873 0.897 0.776 0.780 0.864
End2End (Kolitsas et al., 2018) 0.891 0.933 0.894 0.892 0.760 0.740 0.851
GENRE (Cao et al., 2021) 0.933 0.943 0.899 0.901 0.773 0.874 0.887
ReFinED (Ayoola et al., 2022b) 0.939 0.941 0.918 0.908 0.794 0.874 0.894
ExtEnD (Barba et al., 2022a) 0.926 0.947 0.916 0.918 0.777 0.888 0.895
CoherentED (Xiao et al., 2023b) 0.894 0.963 0.946 0.934 0.811 0.906 0.909
AIDA GENRE (Cao et al., 2021) 0.886 0.881 0.771 0.823 0.719 0.717 0.795
ner4el (Tedeschi et al., 2021) 0.925 0.892 0.695 0.913 0.685 0.640 0.792
ExtEnd (Barba et al., 2022a) 0.900 0.945 0.879 0.889 0.766 0.767 0.858
@@ 4.4.1. Entity Disambiguation (RQ1 and RQ2)
Table 3. Test micro-$F_{1}$ scores on five ED benchmarks. We compare EntGPT-P and EntGPT-I with other open-sourced ED methods. Highest micro-$F_{1}$ scores are highlighted in bold, second highest micro-$F_{1}$ scores are underlined. The experiments are re-computed from source code to compare entity names for evaluation.
@@
We further evaluated our approach on five entity benchmarks from General Entity Annotator Benchmarking Framework (GERBIL) (Usbeck et al., 2015). Because previous work does not evaluate on these datasets, we could only consider open-source entity disambiguation models for comparison. We found that both EntGPT-I and EntGPT-P achieved improvement over the other SOTA models as shown in Table 3. EntGPT-I and EntGPT-P produce an average micro-$F_{1}$ improvement of +10.0%, and +6.3% respectively over the currently highest performing approach (ReFinED). We take special note of the performance gain of the zero-shot EntGPT-P model, which highlights the effectiveness of the multi-step prompting design.
Table 4. Accuracy on six QA datasets. Most accurate zero-shot scores are highlighted in bold, second most accurate zero-shot scores are underlined. Supervised learning methods which are fine-tuned on each dataset individually. † indicates that the results are obtained directly from the source paper.
@@
supervised ROBERTa (Liu et al., 2019) 0.430 0.649 0.827 0.805 0.794 0.759
QAGNN (Yasunaga et al., 2021) 0.444 0.678 0.830 0.734 0.796 0.757
GreaseLM (Zhang et al., 2022a) 0.447 0.669 0.833 0.742 0.796 0.833
DRAGON (Yasunaga et al., 2022) 0.486 0.720 0.840 0.760 0.811 0.768
zero-shot LLAMA-65B (Touvron et al., 2023a) 0.560 0.602 - - 0.828 0.523
FLAN-137B (Wei et al., 2021) 0.631 0.784 - - 0.805 -
Llama-2-70B (Touvron et al., 2023b) 0.574 0.602 - - 0.828 0.507
@@ 4.4.2. Question Answering (RQ1 and RQ2)
4.5. Ablation Study (RQ3)
Table 5. Ablation study on 8 datasets with GPT3.5 backbone. The highest scores are highlighted in bold.
@@ 4.5. Ablation Study (RQ3)
Table 6. Comparison of the ground truth label and the label predicted by the EntGPT-P model. In many error cases the EntGPT-P model produces labels that, in our opinion, are more accurate than the labeled ground truth. In other cases where the ground truth was indeed correct, the errant prediction was close.
@@
ACE04
high
Ministry of Defense (Iran)
Ministry of Defense (Japan)
Step 3
ACE04
low
President of Egypt
President
Step 3
ACE04
low
Gaza City
Gaza Strip
Step 2
ACE04
low
Volvo
Volvo Cars
Step 3
KORE
low
First Ladies of Argentina
First Lady
Step 2
KORE
high
Justin Bieber
Justin I
Step 2
KORE
high
Lady Gaga
Gwen Stefani
Step 3
KORE
high
Paul Allen
NULL
Step 2
AQU
high
Cancer
Lung Cancer
Step 3
AQU
high
Tissue (biology)
Facial tissue
Step 3
CWEB
high
Head
Head (company)
Step 2
CWEB
none
Hillsborough County, Florida
Hillsborough, North Carolina
GT is incorrect
CWEB
low
Lake Wylie
Lake Wylie, South Carolina
Step 3
CWEB
none
Australia Cricket Team
Australia
GT is incorrect
MSN
none
New York City
New York
GT is incorrect
MSN
none
University of Alabama
Alabama Crimson Tide football
GT is incorrect
MSN
high
World Trade Center
Collapse of the World Trade Center
Step 3
OKE15
none
Fellow
Research Fellow
GT is incorrect
OKE15
low
Cambridge
University of Cambridge
Step 2
OKE15
none
Principal (academia)
Head teacher
GT is incorrect
OKE15
low
Faculty (academic staff)
Professor
Step 2
OKE15
none
Officer
Officer (armed forces)
GT is incorrect
OKE16
none
Director (business)
Executive director
GT is incorrect
OKE16
none
Germany
Nazi Germany
GT is incorrect
OKE16
none
Czechs
Czech Republic
GT is incorrect
OKE16
none
Sorbonne
University of Paris
GT is incorrect
REU
none
Georgia Power
Georgia (U.S. state)
GT is incorrect
REU
low
Lloyds Bank of Canada
Lloyds Bank
Step 3
RSS
none
Steve Jobs
Apple Inc.
GT is incorrect
RSS
low
Pro Bowl
Super Bowl
Step 3
RSS
none
Eric Kearney
Cincinnati
GT is incorrect
RSS
high
Cleveland Browns
Cleveland
Step 3
5. A Case Study on Entity Disambiguation
@@
Figure 2. Examples from the entity disambiguation task. We compared GPT3.5, EntGPT-P and EntGPT-I on the OKE15 dataset. Mentions are highlighted within the brackets.
Three example cases from the OKE15 dataset are compared across three models (GPT3.5, EntGPT-P, and EntGPT-I) on disambiguation task are illustrated in Fig. 2. Given the task of linking the mention “Farnborough”, with context provided, EntGPT-I correctly identified “Farnborough, Hampshire”. Yet there were also instances where EntGPT-I failed to correctly link the mention to the desired entity.
6. Case Study on Question Answering (RQ4)
To better understand the results entity linking provides to LLMs, we also conducted a brief error analysis on two of the QA datasets: ARC-C and OBQA. ARC-C is comprised of science exam questions collected from various sources; OBQA is a collection of multiple-choice elementary-level science questions. As reported in Table 4, we performed our experiments with three different models: Llama2-chat, GPT3.5, and EntGPT-I with a GPT3.5 foundation model. We highlight some of the interesting cases as observed in Fig. 3 to assess how different models behave under varying conditions.
@@ 6. Case Study on Question Answering (RQ4)
Figure 3. Examples from two QA tasks: ARC-C (left), and OBQA (right), showing the performance of different language models on a particular question. Here, GPT refers to the foundational GPT3.5 model, and Llama2 refers to the Llama2-70B-chat.
Moreover, we performed an analysis on cases where the questions required both commonsense and logical reasoning. We found that entity-based knowledge provided the models with a better understanding of the context and therefore helps to mitigate hallucination, with some exceptions. As highlighted in Fig 3, the examples that require both commonsense knowledge and logical deduction such as the predators-prey relationship and the shark example do benefit from grounding the entities. However, we did sometimes encounter instances when entity linking did not improve performance. For example, in the soil case, the entity-linking lead the model astray, finding “roots” rather than the more-direct correct answer “microscope”. Although these errors are unfortunate, they still leave the door open for additional improvement in future work.
7. Discussion
In this work, we described the Entity GPT model and two variants: prompt-based (EntGPT-P) and instruction-tuned (EntGPT-I). By linking natural language mentions of concrete entities generated by large language models to their corresponding entities in some knowledge base, we showed that the EntGPT approach reduces hallucination. Furthermore, we also conclude from our extensive experiments that enhancing entity-based knowledge not only helps in entity disambiguation task but also provides better performance in question answering tasks.
This study also motivates us to explore two different follow-up experiments as future work. We first wish to extend the EntGPT framework to perform entity linking (not just entity disambiguation). Compared to entity disambiguation, entity linking not only requires a model to select correct entities but also to extract mentions. Although previous studies (Cao et al., 2021; Xiao et al., 2023a) have explored the application of generative models to entity linking, they do not consistently demonstrate advantages over supervised methods. Second, we wish to dive deeper into how and why entity disambiguation improves question answering performance. These initial observations point to entity-correlation as a crucial factor leading to the better performance, so it may be wise to incorporate entity correlation more deeply into the question answering paradigm.

### v3

@@ 4.2 Comparison Models
  • 1.
    End2End Kolitsas-CONLL'18-end2end: The End-to-End Neural Entity Linking (End2End) is a global ED and entity linking system based on Deep-ED Ganea-EMNLP'17-deep_ed, but with a more robust recurrent neural network (RNN) architecture that considers mention-context and mention-mention interactions.
  • 2.
    REL Hulst-SIGIR'20-REL: The Radboud Entity Linker (REL) is an ED and entity linking package which combines the NER method FLAIR Akbik-NAACL'19-FLAIR and ED method mulrel-nel Le-ACL'18-mulrel-nel.
  • 3.
    GENRE DeCao-ICLR'21-GENDRE: The Generative Entity REtrieval model (GENRE) is a generation-based model that re-casts the ED task as an entity name generation process.
  • 4.
    ner4el Tedeschi-EMNLP'21-NER4EL: The Named Entity Recognition for Entity Linking (NER4EL) model uses named entity recognition to perform the entity linking task, which is similar to the entity disambiguation task.
  • 5.
    ReFinED Ayoola-NAACL'22-ReFinED: The Representation and Fine-grained typing for Entity Disambiguation (ReFinED) model is an efficient end-to-end zero-shot-capable EL model is a recent system that considers Wikipedia entries as extra features to the ED and entity linking tasks.
  • 6.
    ExtEnd barba2022extend: The Extractive Entity Disambiguation (ExtEnd) model is a transformer-based framework where the disambiguation task is recast as a machine reading comprehension task.
  • 7.
    CoherentED Xiao-EMNLP'23-CoherentED: The Coherent Entity Disambiguation (CoherentED) model aims to enhance the coherence of entity predictions and achieves state-of-the-art performance on most datasets.
@@
    ROBERTa liu2019roberta: The Robustly Optimized BERT Pretraining Approach (ROBERTa) is an optimized version of the Bidirectional Encoder Representations from Transformers (BERT) devlin2018bert model, which uses Masked Language Modeling (MLM) as its pre-training objective. The pre-trained model is fine-tuned on each question answering benchmark in a fully supervised manner.
  • 2.
@@
    QAGNN & GreaseLM yasunaga2021qa; zhang2022greaselm: The Question Answering Graph Neural Network (QAGNN) and the Graph REASoning Enhanced Language Model (GreaseLM) both fine-tune an LLM using a knowledge graph (KG) on downstream tasks. These architectures are broadly referred to as KG-augmented question answering models.
  • 3.
@@
    DRAGON yasunaga2022deep: Deep Bidirectional Language-Knowledge Graph Pretraining (DRAGON) uses a knowledge base to enhance the pre-training stage and also performs fine-tuning on downstream tasks in a supervised manner.
@@
In the zero-shot setting, we use the following LLMs as foundation models: Large Language Model Meta AI (Llama, Llama2, & Llama2-chattouvron2023llama; Touvron-arXiv'23-Llama2, the Fine-tuned LAnguage Net (FLAN-137B) wei2021finetuned, and the Generative Pre-trained Transformer (GPT3.5) Ouyang-NIPS'22-InstructGPT.
Table 1: Statistics for 11 experimental datasets.
# of Docs # of Mentions
AIDA 231 4485
MSN 20 656
AQU 50 727
ACE04 35 257
CWEB 320 11154
WIKI 319 6793
KORE 50 144
OKE15 101 536
OKE16 173 288
REU 113 650
RSS 357 524
Table 2: Examples from the entity disambiguation task. We compared GPT3.5, EntGPT-P and EntGPT-I on the OKE15 dataset. Mentions are underlined with blue color.
{tabularx}
@ p.52 X @ Example Model Prediction
Blackett spent ten years working at the Cavendish Laboratory as an experimental physicist with Ernest Rutherford and in 1923 became a fellow of King’s college, Cambridge.
GPT3.5: IEEE
EntGPT-P: None
EntGPT-I: Fellow
The man spent time at the Royal Aircraft Establishment in Farnborough, where he made a major contribution to the design of the Mark XIV bomb sight.
GPT3.5: Royal Aircraft Establishment
EntGPT-P: Royal Aircraft Establishment
EntGPT-I: Farnborough, Hampshire
Here Lutuli was confirmed in the Methodist Church and became a lay preacher.
GPT3.5: Methodist Church of Great Britain
EntGPT-P: United Methodist Church
EntGPT-I: Methodism
Table 3: Examples from two QA tasks: ARC-C (left), and OBQA (right), showing the performance of different language models on a particular question. Here, GPT3.5 refers to the foundational GPT3.5 model, and LLAMA2 refers to the Llama2-70B-chat.
Example Model Prediction
ARC-C An astronaut drops a 1.0 kg object and a 5.0 kg object on the Moon. Both objects fall a total distance of 2.0 m vertically. Which of the following best describes the objects after they have fallen a distance of 1.0 m. (A) They have each lost kinetic energy.
(B) They have each gained the same amount of potential energy.
(C) They have each lost the same amount of potential energy.
(D) They have each gained one-half of their maximum kinetic energy
GPT3.5: B EntGPT-I: C LLAMA2: A
In the equatorial Pacific Ocean, upwelling moves cold, deep sea water to the surface. This water is rich in nutrients and dissolved carbon dioxide gas. During an El Niño event, upwelling declines. This causes surface waters to rise in temperature. Which of the following is most likely to occur in the equatorial Pacific Ocean during an El Niño year? (A) decrease in release of carbon dioxide to the atmosphere
(B) increase in photosynthesis by algae in surface waters
(C) increase in dissolved oxygen in surface waters
(D) decrease in precipitation over ocean water
GPT3.5: B EntGPT-I: A LLAMA2: B
OBQA Predators eat:
(A) lions (B) humans (C) bunnies (D) grass
GPT3.5: B EntGPT-I: C LLAMA2: A
The summer solstice in the northern hemisphere is four months before:
(A) May (B) July (C) April (D) October
GPT3.5: B EntGPT-I: B LLAMA2: C
They studied the soil by using:
(A) plants (B) a telescope (C) roots (D) a microscope
GPT3.5: D EntGPT-I: C LLAMA2: D
If you were attacked by a shark and had to punch it sharply where it pulls in air from, you’d use your hand to make contact with: (A) its snout (B) its gills (C) its nose (D) its belly GPT3.5: A EntGPT-I: B LLAMA2: A
4.3 Evaluation Metrics
@@
To maintain a fair comparison across datasets, we use the in-KB micro-F1 score as our evaluation metric for ED following the example of Guo and Barbosa Guo-Semantic_Web'16-Random_Walk. Specifically, being in-KB requires that ground truth mentions correspond to existing knowledge base entries. Empty or out-of-KB entities are removed in the evaluation process. The mean Micro-F1 score is taken per-mention. Although a model may predict non-entities, each mention will have some corresponding ground truth entity. For question answering, we use accuracy as the evaluation metric.
4.4 Results and Findings
@@
4.4.1 Entity Disambiguation (RQ1 and RQ2)
Table 4: Test micro-F1 scores on standard six benchmarks. With each setting, best scores are highlighted in bold. † indicates that the results are obtained directly from the source paper.
Model AIDA MSN AQU ACE04 CWEB WIKI avg
Wiki+AIDA REL Hulst-SIGIR'20-REL 0.928 0.935 0.873 0.897 0.776 0.780 0.864
End2End Kolitsas-CONLL'18-end2end 0.891 0.933 0.894 0.892 0.760 0.740 0.851
GENRE DeCao-ICLR'21-GENDRE 0.933 0.943 0.899 0.901 0.773 0.874 0.887
ReFinED Ayoola-NAACL'22-ReFinED 0.939 0.941 0.918 0.908 0.794 0.874 0.894
ExtEnD Barba-ACL'22-Extend 0.926 0.947 0.916 0.918 0.777 0.888 0.895
CoherentED Xiao-EMNLP'23-CoherentED 0.894 0.963 0.946 0.934 0.811 0.906 0.909
AIDA GENRE DeCao-ICLR'21-GENDRE 0.886 0.881 0.771 0.823 0.719 0.717 0.795
ner4el Tedeschi-EMNLP'21-NER4EL 0.925 0.892 0.695 0.913 0.685 0.640 0.792
ExtEnd Barba-ACL'22-Extend 0.900 0.945 0.879 0.889 0.766 0.767 0.858
@@ 4.4.1 Entity Disambiguation (RQ1 and RQ2)
Table 5: Test micro-$F_{1}$ scores on five ED benchmarks. We compare EntGPT-P and EntGPT-I with other open-sourced ED methods. Highest micro-$F_{1}$ scores are highlighted in bold, second highest micro-$F_{1}$ scores are underlined. The experiments are re-computed from source code to compare entity names for evaluation.
@@
We further evaluated our approach on five entity benchmarks from General Entity Annotator Benchmarking Framework (GERBIL) Usbeck-WWW'15'-GERBIL. Because previous work does not evaluate on these datasets, we could only consider open-source entity disambiguation models for comparison. We found that both EntGPT-I and EntGPT-P achieved improvement over the other SOTA models as shown in Table 5. EntGPT-I and EntGPT-P produce an average micro-$F_{1}$ improvement of +10.0%, and +6.3% respectively over the currently highest performing approach (ReFinED). We take special note of the performance gain of the zero-shot EntGPT-P model, which highlights the effectiveness of the multi-step prompting design.
Table 6: Accuracy on six QA datasets. Most accurate zero-shot scores are highlighted in bold, second most accurate zero-shot scores are underlined. Supervised learning methods which are fine-tuned on each dataset individually. † indicates that the results are obtained directly from the source paper.
@@
supervised ROBERTa liu2019roberta 0.430 0.649 0.827 0.805 0.794 0.759
QAGNN yasunaga2021qa 0.444 0.678 0.830 0.734 0.796 0.757
GreaseLM zhang2022greaselm 0.447 0.669 0.833 0.742 0.796 0.833
DRAGON yasunaga2022deep 0.486 0.720 0.840 0.760 0.811 0.768
zero-shot LLAMA-65B touvron2023llama 0.560 0.602 - - 0.828 0.523
FLAN-137B wei2021finetuned 0.631 0.784 - - 0.805 -
Llama-2-70B Touvron-arXiv'23-Llama2 0.574 0.602 - - 0.828 0.507
@@ 4.4.2 Question Answering (RQ1 and RQ2)
4.5 Ablation Study (RQ3)
Table 7: Ablation study on 8 datasets with GPT3.5 backbone. The highest scores are highlighted in bold.
@@ 4.5 Ablation Study (RQ3)
Table 8: Comparison of the ground truth label and the label predicted by the EntGPT-P model. In many error cases the EntGPT-P model produces labels that, in our opinion, are more accurate than the labeled ground truth. In other cases where the ground truth was indeed correct, the errant prediction was close.
@@
ACE04 high Ministry of Defense (Iran) Ministry of Defense (Japan) Step 3
ACE04 low President of Egypt President Step 3
ACE04 low Gaza City Gaza Strip Step 2
ACE04 low Volvo Volvo Cars Step 3
KORE low First Ladies of Argentina First Lady Step 2
KORE high Justin Bieber Justin I Step 2
KORE high Lady Gaga Gwen Stefani Step 3
KORE high Paul Allen NULL Step 2
AQU high Cancer Lung Cancer Step 3
AQU high Tissue (biology) Facial tissue Step 3
CWEB high Head Head (company) Step 2
CWEB none Hillsborough County, Florida Hillsborough, North Carolina GT is incorrect
CWEB low Lake Wylie Lake Wylie, South Carolina Step 3
CWEB none Australia Cricket Team Australia GT is incorrect
MSN none New York City New York GT is incorrect
MSN none University of Alabama Alabama Crimson Tide football GT is incorrect
MSN high World Trade Center Collapse of the World Trade Center Step 3
OKE15 none Fellow Research Fellow GT is incorrect
OKE15 low Cambridge University of Cambridge Step 2
OKE15 none Principal (academia) Head teacher GT is incorrect
OKE15 low Faculty (academic staff) Professor Step 2
OKE15 none Officer Officer (armed forces) GT is incorrect
OKE16 none Director (business) Executive director GT is incorrect
OKE16 none Germany Nazi Germany GT is incorrect
OKE16 none Czechs Czech Republic GT is incorrect
OKE16 none Sorbonne University of Paris GT is incorrect
REU none Georgia Power Georgia (U.S. state) GT is incorrect
REU low Lloyds Bank of Canada Lloyds Bank Step 3
RSS none Steve Jobs Apple Inc. GT is incorrect
RSS low Pro Bowl Super Bowl Step 3
RSS none Eric Kearney Cincinnati GT is incorrect
RSS high Cleveland Browns Cleveland Step 3
5 A Case Study on Entity Disambiguation
@@
Three example cases from the OKE15 dataset are compared across three models (GPT3.5, EntGPT-P, and EntGPT-I) on disambiguation task are illustrated in Fig. LABEL:ed_case. Given the task of linking the mention “Farnborough”, with context provided, EntGPT-I correctly identified “Farnborough, Hampshire”. Yet there were also instances where EntGPT-I failed to correctly link the mention to the desired entity.
6 Case Study on Question Answering (RQ4)
To better understand the results entity linking provides to LLMs, we also conducted a brief error analysis on two of the QA datasets: ARC-C and OBQA. ARC-C is comprised of science exam questions collected from various sources; OBQA is a collection of multiple-choice elementary-level science questions. As reported in Table 6, we performed our experiments with three different models: Llama2-chat, GPT3.5, and EntGPT-I with a GPT3.5 foundation model. We highlight some of the interesting cases as observed in Fig. LABEL:fig:obqa to assess how different models behave under varying conditions.
@@ 6 Case Study on Question Answering (RQ4)
Moreover, we performed an analysis on cases where the questions required both commonsense and logical reasoning. We found that entity-based knowledge provided the models with a better understanding of the context and therefore helps to mitigate hallucination, with some exceptions. As highlighted in Fig LABEL:fig:obqa, the examples that require both commonsense knowledge and logical deduction such as the predators-prey relationship and the shark example do benefit from grounding the entities. However, we did sometimes encounter instances when entity linking did not improve performance. For example, in the soil case, the entity-linking lead the model astray, finding “roots” rather than the more-direct correct answer “microscope”. Although these errors are unfortunate, they still leave the door open for additional improvement in future work.
7 Discussion
In this work, we described the Entity GPT model and two variants: prompt-based (EntGPT-P) and instruction-tuned (EntGPT-I). By linking natural language mentions of concrete entities generated by large language models to their corresponding entities in some knowledge base, we showed that the EntGPT approach exhibit exceptional ability in entity disambiguation. Furthermore, we also conclude from our extensive experiments that enhancing entity-based knowledge not only helps in entity disambiguation task but also provides better performance in question answering tasks.
This study also motivates us to explore two different follow-up experiments as future work. We first wish to extend the EntGPT framework to perform entity linking (not just entity disambiguation). Compared to entity disambiguation, entity linking not only requires a model to select correct entities but also to extract mentions. Although previous studies DeCao-ICLR'21-GENDRE; Xiao-EMNLP'23-Instruction have explored the application of generative models to entity linking, they do not consistently demonstrate advantages over supervised methods. Second, we wish to dive deeper into how and why entity disambiguation improves question answering performance. These initial observations point to entity-correlation as a crucial factor leading to the better performance, so it may be wise to incorporate entity correlation more deeply into the question answering paradigm.